Many scientists believe that large asteroid struck the earth that kill off dinosaurs. Fragment of asteroid had found out and contain 86% of its original uranium-238. How old is the fragment of the asteroid? The half-life of uranium-238 is 4.5 billion years

1
Expert's answer
2021-12-21T04:19:25-0500

Decay equation


"N(t)=N_0e^{-\lambda t}"

The half-life of uranium-238 is 4.5 billion years


"\dfrac{N_0}{2}=N_0e^{-\lambda (4.5\times10^9)}"

"e^{\lambda (4.5\times10^9)}=2"

"\lambda=\dfrac{\ln 2}{4.5\times10^9}"

Then


"N(t)=N_0e^{-(\ln2/(4.5\times10^9))t}"

"N(t)=N_0(2)^{-t/(4.5\times 10^9)}"

Given "N(t_1)=0.86N_0"


"0.86N_0=N_0(2)^{-t/(4.5\times 10^9)}"

"t_1=-4.5\times10^9\cdot\big(\dfrac{\ln 0.86}{\ln2}\big)"

"t_1=979\times10^6 \ years"

The fragment of the asteroid is 979 million years old.